The Chinese HOWTO

Hong Zhang, hongz@seas.ucla.edu v0.1, 2 April 1997 This is the HOWTO document for using Chinese in Linux environment. It describes how to obtain, install, and configure various Chinese software. The most up-to-date version of this Chinese HOWTO may be found at

turb10.seas.ucla.edu:/pub/

Contents

1 Introduction 3

2 The Chinese X Windows 3

2.1 Where to get it . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.1 CXwin 3.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.2 CXwin 3.2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.3 Patches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.2 How to install it . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 The CXterm and Hztty 5 3.1 Getting the cxterm package and hztty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Installing cxterm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Installing hztty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.4 Color patch for cxterm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.5 Trouble shooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Chinese Editors 7

4.1 Celvis and CVim - clones of vi/ex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.1.1 Getting the celvis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.1.2 Installing celvis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.1.3 Cvim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.2 Cjoe - Joe’s Own Chinese Editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2.1 Getting cjoe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2.2 Compiling and Installing cjoe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.3 Cemacs and CChelp For Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.3.1 Getting Cemacs and CChelp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.3.2 Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.4 Mule and Xemacs 20.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.4.1 Getting Mule-2.3 and the Patch for Linux . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.4.2 Compile and Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.4.3 Use Chinese in Mule-2.3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4.4 Xemacs 20.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5 Installing Chinese fonts 10

5.1 Installing X window fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.1.1 Where to get GB and BIG5 fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.1.2 Installing X fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

5.2 HBF fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6 Printing Utilities 12

6.1 CNPRINT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.1.1 Where is cnprint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.1.2 Install cnprint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6.2 GB2PS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.2.1 Where to find GB2PS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.2.2 Installing GB2PS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

6.3 DTop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

7 Displaying Chinese in Netscape 13

8 Chinese Tex 14 8.1 CJK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.2 Poor Man’s Tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.3 ChTex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

9 Using Chinese in Console – Chdrv and WZCE 14 9.1 Getting chdrv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 9.2 Installing chdrv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 9.3 GB Chinese Console - WZCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

9.3.1 Where is it . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

9.3.2 Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

10 Chinese WordProcessor – ChinesePower 15

10.1 Where to get chpower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

10.2 Installing chpower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

11 Wabi 2.2 and Chinese window system 16

12 Other Cool Stuffs 16

12.1 Crxvt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

12.1.1 Getting Crxvt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

12.1.2 Compiling and Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

12.2 Ktty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

12.3 TinTin++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

13 Need more help ? 17

14 Appendix - Chinese Software Catalogue of FTP.IFCSS.ORG 17

1 Introduction

This is an overdue document for using Chinese in Linux. I have been expecting this HOWTO for a long time. Obviously there is no such HOWTO available on Internet, I decided to write one by myself. It is nothing else but an attempt to make life a little bit easer. Please send your comments and suggestions to hongz@seas.ucla.edu. There is no copyright issue associated with this document. You may do whatever you want with this HOWTO except for suing me for the damage caused by this document or bringing me into a lawsuit.

You can get most of the Chinese Software at:

     <http://cnapps.ifcss.org> and  <ftp://linux.csie.nctu.edu.tw:/pub/linux>

2 The Chinese X Windows

The X Window system is a large and powerful graphics environment for UNIX. XFree86 is a free port of the MIT X Window System (X11R6). The Chinese X Windows is a patch to the XFree86 for displaying chinese charactors in X Window environment. The author is Shin-Ray Lee (srlee@csie.nctu.edu.tw). Current version of CXwin only supports BIG5 coding. It is available for both XFree86 3.1 and XFree86 3.2 . CXwin gives your Chinese popup menu, Chinese window title for various window managers and also can display Chinese characters in some applications.

2.1 Where to get it

2.1.1 CXwin 3.1

The XFree86 3.1 version of Chinese X Windows can be found at the major chinese software ftp site

ftp.ifcss.org:/pub/software/linux/X11R6/CXwin/elf or its mirror sites. You only need to get one X

server for yourself. In my case, I need XF86 SVGA.gz. But XF86 VGA16.gz should work for most of the graphics cards.

2.1.2 CXwin 3.2

As I know, the latest version for CXwin is XFree86 3.2 . You can fetch the linux ELF format of CXwin 3.2

at linux.csie.nctu.edu.tw:/pub/linux/CXwin/elf

2.1.3 Patches

If you have the source for XFree86, you can get patch at the same place and compile the binary file by yourself.

2.2 How to install it

Uncompress the gzipped file, in my case it is /tmp/XF86 SVGA.gz

     # gzip -d XF86_SVGA.gz

Become superuser (root) and make a backup copy of your orignal X server

     # cd /usr/X11R6/bin
     # mv XF86_SVGA XF86_SVGA_BACKUP

Move the CXwin to its place

     # mv /tmp/XF86_SVGA /usr/X11R6/bin
     # chown root.bin XF86_SVGA
     # chmod 4755 XF86_SVGA

CXwin also needs four BIG5 fonts to work properly, namely taipei15, taipei16, taipeik20 and taipeik24. Please refer to “Section 5” for how to intall Chinese fonts.

CXwin 3.1 is buggy and unstable. It is recommended that you upgrade to XFree86 3.2 and use CXwin 3.2. CXwin is in alpha state, you may use it at your own risk. But I have been using it for over three months, it works just great. At <http://turb10.seas.ucla.edu/~ hongz>, you will find some nice screenshots of my desktop.

3 The CXterm and Hztty

CXterm is a terminal simulator just like xterm but with the ability of displaying and inputting chinese charactors. It supports GB, BIG5 coding. Along with hztty, you can also read HZ encoded text which is commonly used in chinese newsgroup. You may read CXTERM.FAQ (cs.purdue.edu:ftp/pub/ygz) for installation details. Yongguang Zhang is the author of this great software.

3.1 Getting the cxterm package and hztty

cxterm5.0.p3.tar.gz is the latest version of cxterm (version 5.0). This single package includes cxterm and chinese fonts. It is very easy to install the cxterm. You can get cxterm and hztty from

cxterm5.0.p3.tar.gz    ftp.ifcss.org:/pub/software/x-win/cxterm
hztty-2.0.tar.gz       ftp.ifcss.org:/pub/software/unix/convert

3.2 Installing cxterm

Unpacking the cxterm package

     # tar -xvzf cxterm5.0.p3.tar.gz

Which will creat a new directory cxterm-5.0

     # cd cxterm-5.0
     # ./config.sh

If you want to install cxterm for all the users on the system, you need to become root before you run the command “./config.sh”. Now you can follow the instruction to compile and to install cxterm.

-----------------------------------------------------------------------------
    --- BASIC MENU ---
  0. Read COPYRIGHT Notice
  1. Compile, Install, and Configure "CXTERM 5.0" in One Step
    --- OPTION MENU ---
  2. Compile cxterm (not to install)
  3. Install cxterm (after successful compilation in 2)
  4. Install additional Chinese font(s) for your X window
  5. Configure your account for using cxterm (after installation in 3)
  x. Exit
-----------------------------------------------------------------------------
Please choose (0/1/2/3/4/5/x) :

Choose option 1 if you want the installation script to do everything for you. At the middle, you will be prompt to enter the installing directory. In my case, I installed cxterm in /usr/local/chinese. There are also two fonts come with the cxterm package, choose option 1 and 3 will install the fonts automatically. You can also install extra fonts using option 4. After installing cxterm, make sure the executables cxterm and CXterm in your search path. For C shell and Tcsh

     # set path (/usr/local/chinese/bin $path)

For sh and bash

     # setenv PATH=$PATH:/usr/local/chinese/bin ; export PATH

CXterm is a shell script for loading X window resources and bringing up the cxterm. For using cxterm in GB mode, try this

     # CXterm -gb

Or in BIG5 mode

     # CXterm -big5

3.3 Installing hztty

Hztty translates Chinese between different encoding. Unpacking hztty-2.0.tar.gz

     # tar -xvzf hztty-2.0.tar.gz
     # cd hztty-2.0
     # make linux

After compiling, move the executable hztty to your bin directory and the man page to your man directory.

     # mv hztty /usr/local/bin
     # chmod 555 /usr/local/bin/hztty
     # cp hztty.1 /usr/local/man/man1
     # chmod 444 /usr/local/man/man1/hztty.1

Read the hztty man page for how to use hztty.

3.4 Color patch for cxterm

There are also one patch file cxterm-5.0.p3-color.patch.gz locate in the same directory as

cxterm5.0.p3.tar.gz at ftp.ifcss.org:/pub/software/x-win/cxterm. You can patch the cxterm

source to get ANSI color for cxterm. Here suppose you have the cxterm source tree in /tmp/cxterm-5.0

     # cp cxterm-5.0.p3-color.patch.gz /tmp
     # gzip -d cxterm-5.0.p3-color.patch.gz
     # patch < cxterm-5.0.p3-color.patch
     # cd cxterm-5.0
     # ./config.sh

The color-cxterm is kind of cool. But There is one annoying thing when I try to input Chinese in color-cxterm: the input area at the low bottom of cxterm does not redraw properly, I have to press “Ctrl-L” to redraw the whole screen. This maybe just a termcap problem. Hope someone will give me a hint.

3.5 Trouble shooting

Sometimes Cut & Paste may not working in cxterm shell prompt. Tcsh strips the 8th bit. You need to type “setenv LANG” to correct this. Also try “stty pass8”.

4 Chinese Editors

4.1 Celvis and CVim - clones of vi/ex

C-Elvis is a clone of vi/ex, the standard UNIX editor. C-Elvis supports editing of Chinese text, English text, and the mixture of both. C-Elvis supports nearly all the vi/ex commands, in both visual mode and ex (colon) mode, for both Chinese and English characters. You can use C-Elvis for both GB2312-80 standard and Big5 encoding Chinese text. You need to use it in a Chinese terminal such as cxterm to input/output Chinese character.

4.1.1 Getting the celvis

The ftp site for celvis is

celvis-1.3.tar.gz               ifcss.org:/pub/software/unix/editor

4.1.2 Installing celvis

     # tar -xvzf celvis-1.3.tar.gz
     # cd celvis

You need to edit tmp.c, delete line 93-95

#if OS9
                  if we don't have write permission...
#endif

Then you can compile it by

     # cp Makefile.s5 Makefile
     # make install

Because the Makefile is not for Linux, you may get a lot of warning messages. But you can ignore all of them. The celvis will installed in in /usr/local/bin.

4.1.3 Cvim

CVim is a patch to add Chinese support for vim-4.2. CVim includes some vi’s features which are not supported by celvis-1.3, such as line number, line wrap, and large file editing. you can get

vim-4.2-Chinese-patch and vim-4.2.tar.gz at ftp.csie.nctu.edu.tw:/pub/Unix/Chinese/cvim
     # tar -xvzf vim-4.2.tar.gz
     # cd vim-4.2/src
     # patch < ../../vim-4.2-Chinese-patch

The file “vim-4.2/src/feature.h” can be edited to match your preferences. Compiling is simple, just do

     # make
     # make install

4.2 Cjoe - Joe’s Own Chinese Editor

JOE is a professional freeware ASCII text screen editor for UNIX. JOE has the feel of most IBM PC text editors. Although I am not using it myself, it is definitly a great editor.

4.2.1 Getting cjoe

You can find cjoe-2.8.tgz at the same place as celvis-1.3.tar.gz
cjoe-2.8.tgz            ifcss.org:/pub/software/unix/editor

4.2.2 Compiling and Installing cjoe

Edit the Makefile file if you want to change the locations for binaries, keymap initialization files and man page. Otherwise, just do the following

     # make
     # make install

4.3 Cemacs and CChelp For Emacs

Cemacs is a method of using GNU Emacs to display and edit Chinese files. To use Cemacs, it is necessary to run Emacs on a Chinese terminal or under a Chinese terminal emulator program such as cxterm. CCHELP is a system to obtain instant helpful information on Chinese characters. Once CCHELP is installed, you can click the mouse on any Chinese character and a help messag will appear giving the character itself, its PinYin pronunciation, and an indication of its English meaning. Both GB- and Big5-coded Chinese are supported.

4.3.1 Getting Cemacs and CChelp

Both cemacs and cchelp can be fetched at ftp.math.psu.edu:/pub/simpson/chinese

4.3.2 Installing

Follow the README files for cemacs and cchelp.

4.4 Mule and Xemacs 20.0

4.4.1 Getting Mule-2.3 and the Patch for Linux

mule-2.3.tar.gz            ftp.ifcss.org:/pub/software/mule/editor
                           sunsite.unc.edu:/pub/Linux/distribution/je/sources/mule
mule-2.3+lx.diff.gz        sunsite.unc.edu:/pub/Linux/distribution/je/sources/mule

4.4.2 Compile and Installing

Unpack the package and apply the patch

     # tar -xvzf mule-2.3.tar.gz
     # patch < mule-2.3+lx.diff
     # cd mule-2.3/
     # ./configure "i386-*-linux" --with-x11 --with-x-toolkit --with-gcc

If you want to customize your mule, read the file “INSTALL” and try “./configure --help”. After the configuration, you need to modify the file “src/Makefile” a little bit, change “-lcurses” to “-lncurses”. Now

     # make
     # make install

The default installation directory is /usr/local.

5. Installing Chinese fonts 10

4.4.3 Use Chinese in Mule-2.3

If you have fonts intalled, you can use mule for input/output Chinese. Most of the Chinese fonts are 16 pts or 24 pts, so try

# mule -fn 8x16 &

Or

# mule -fn 12x24 &

In mule type M-x load-library RETURN chinese RETURN . The command “Ctrl-]” toggles input methods.

4.4.4 Xemacs 20.0

Haven’t try yet.

5 Installing Chinese fonts

5.1 Installing X window fonts

5.1.1 Where to get GB and BIG5 fonts

Most likely you may already have some Chinese fonts sitting in your system. You can find GB fonts by

# xlsfonts | grep gb
-cclib-song-medium-r-normal--0-0-72-72-c-0-gb2312.1980-1
-cclib-song-medium-r-normal--16-160-72-72-c-160-gb2312.1980-1
-isas-fangsong ti-medium-r-normal--0-0-72-72-c-0-gb2312.1980-0
-isas-fangsong ti-medium-r-normal--16-160-72-72-c-160-gb2312.1980-0
-isas-song ti-medium-r-normal--0-0-72-72-c-0-gb2312.1980-0
-isas-song ti-medium-r-normal--16-160-72-72-c-160-gb2312.1980-0
-isas-song ti-medium-r-normal--24-240-72-72-c-240-gb2312.1980-0
hanzigb13fs
hanzigb13st
hanzigb16fs
hanzigb16st
hanzigb24st

The last five fonts in this case are just the aliases of the first five fonts. Or to find BIG5 fonts

# xlsfonts | grep big5
-kc-kai-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-kc-kai-medium-r-normal-fantizi-20-200-75-75-c-200-big5.eten.3.10-1
-kc-kai-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1

5. Installing Chinese fonts 11

-kc-ming-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-kc-ming-medium-r-normal-fantizi-15-150-75-75-c-160-big5.eten.3.10-1
-kc-ming-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1
-uw-ming-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-uw-ming-medium-r-normal-fantizi-16-160-75-75-c-160-big5.eten.3.10-1
-uw-ming-medium-r-normal-fantizi-20-200-75-75-c-200-big5.eten.3.10-1
-uw-songti-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-uw-songti-medium-r-normal-fantizi-20-200-75-75-c-200-big5.eten.3.10-1
-uw-songti-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1

If you don’t have any Chinese fonts installed or just want more fonts, you may look into

ftp.ifcss.org:/pub/software/fonts/ where gb/bdf/ is for GB fonts, big5/bdf is for BIG5 fonts.

5.1.2 Installing X fonts

Suppose you get BIG5 font eb5-24k2.bdf.gz and GB font gb24st.bdf.gz. To build the fonts into .pcf format

# gzip -cd eb5-24k2.bdf.gz | bdftopcf -t > eb5-24k2.pcf
# gzip -cd gb24st.bdf.gz | bdftopcf -t > gb24st.pcf
# compress *.pcf
Install the fonts into /usr/X11R6/lib/X11/fonts/misc
# mv *.pcf.Z /usr/X11R6/lib/X11/fonts/misc/
# cd /usr/X11R6/lib/X11/fonts/misc
# mkfontdir .

After installing the fonts, looking into the file fonts.dir, which now contains two lines

eb5-24k2.pcf.Z -kc-kai-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1
gb24st.pcf.Z -isas-song ti-medium-r-normal--24-240-72-72-c-240-gb2312.1980-0

You may need to make aliases for these fonts, which will save you a lot of typing when bringing up Chinese software, add two lines to the file fonts.alias

gb24st    "-isas-song ti-medium-r-normal--24-240-72-72-c-240-gb2312.1980-0"
taipeik24 "-kc-kai-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1"

Now inform your X window system server by

# xset fp rehash

If everything goes well, you can use your new fonts now

# cxterm -GB -fn 12x24 -fh gb24st &

5.2 HBF fonts

HBF fonts contain header files and bitmap font files. You need to get both header file and font files to

install one HBF font. HBF fonts are in ftp.ifcss.org:/pub/software/fonts .

Suppose you want to use font ccfs24.hbf, a GB, FangSong, Jianti font, you should get three files ccfs24.hbf, cclib.n24 and ccsys.24. The file 00index.txt in the ifcss fonts directory gives you filenames of each HBF font.

6 Printing Utilities

6.1 CNPRINT

Cnprint is an utility to print Chinese text or convert Chinese text to PostScript. It works just like a print command and support GB, HZ and BIG5 formats.

6.1.1 Where is cnprint

cnprint260.tar.gz can be found at ftp.ifcss.org:/pub/software/unix/print

6.1.2 Install cnprint

The package cnprint260.tar.gz contains only four files

     # ls
     cnprint.1     cnprint.c     cnprint.cmd   cnprint.help

You can compile it by

     # gcc cnprint.c -o cnprint
     # mv cnprint /usr/local/bin
     # mv cnprint.1 /usr/local/man/man1

For cnprint to work properly, you need to specify the full path for the fonts files (HBF fonts). In csh or

tcsh
     # setenv HBFPATH "/usr/local/chinese/HBF/"
In sh or Bash
     # set HBFPATH="/usr/local/chinese/HBF/"; export HBFPATH

The file cnprint.cmd contains the default settings for cnprint. It should be put in $HBFPATH

     # cp cnprint.cmd $HBFPATH

If you convert Chinese text file to PostScript use the command “cnprint -w FILENAME”

6.2 GB2PS

GB2PS is another printing utility support GB and HZ encoding. It comes with a cover page for printing HXWZ.

6.2.1 Where to find GB2PS

Package:
gb2ps.2.02.tar.gz       ftp.ifcss.org:/pub/software/unix/print
fonts:
csong24.ccf     ckai24.ccf   ftp.ifcss.org:/pub/software/fonts/gb/misc
cfang24.ccf     chei24.ccf

Put fonts in some place. Mine is in /usr/local/chinese/CFONT

6.2.2 Installing GB2PS

Before you compile gb2ps, you need to change two settings in the Makefile, the fonts directory and cover page location

     CFONT=/usr/local/chinese/CFONT/
     COVERPAGE=/usr/local/chinese/lib/cover.ps

Then just type

     # make
     # cp gb2ps /usr/local/bin

6.3 DTop

Don’t know much about it

7 Displaying Chinese in Netscape

There are many web pages on the Internet devoting to this topic. Netscape (better use version 2.1 and up) supports both GB and BIG5. Hare are three steps to view Chinese(BIG5) in Netscape

1. Setup X fonts as in “Section 5”

  2. Bring up netscape and choose menu Options/Document Encoding/Tranditional Chinese (Big5)

3. In menu Options/General Preferences/fonts choose the font you want to use.

8. Chinese Tex 14

But you may wonder why you can only use Fixed(HKU) Size 16.0. It turns out you need to edit the /usr/X11R6/lib/X11/app-defaults/Netscape to get more fonts. Suppose you have the following BIG5 fonts

# xlsfonts | grep big5
-hku-fixed-medium-r-normal--0-0-72-72-c-0-big5.hku-0
-hku-fixed-medium-r-normal--16-160-72-72-c-160-big5.hku-0
-kc-kai-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-kc-kai-medium-r-normal-fantizi-20-200-75-75-c-200-big5.eten.3.10-1
-kc-kai-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1
-kc-ming-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-kc-ming-medium-r-normal-fantizi-15-150-75-75-c-160-big5.eten.3.10-1
-kc-ming-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1
-uw-ming-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-uw-ming-medium-r-normal-fantizi-16-160-75-75-c-160-big5.eten.3.10-1
-uw-ming-medium-r-normal-fantizi-20-200-75-75-c-200-big5.eten.3.10-1
-uw-songti-medium-r-normal-fantizi-0-0-75-75-c-0-big5.eten.3.10-1
-uw-songti-medium-r-normal-fantizi-20-200-75-75-c-200-big5.eten.3.10-1
-uw-songti-medium-r-normal-fantizi-24-240-75-75-c-240-big5.eten.3.10-1
Add this line to your /usr/X11R6/lib/X11/app-defaults/Netscape
*documentFonts.charset*big5.eten.3.10-1:              x-x-big5

Now netscape can use all of them.

8 Chinese Tex

Don’t know much about these packages

8.1 CJK

8.2 Poor Man’s Tex

8.3 ChTex

9 Using Chinese in Console – Chdrv and WZCE

Chdrv is a Chinese Terminal Simulator.It can display Chinese without X-Windows. It bases on VGALIB, and a VGA or super VGA monitor is required. It supports BIG5 coding only, but you can view GB files by

using hztty

9.1 Getting chdrv

chdrvbin-1.0.7.tar.gz      sunsite.unc.edu:/pub/Linux/util/nls
                           ftp.ifcss.org:/pub/software/linux/tty
chdrvfont.tar.gz (HBF)     ftp.ifcss.org:/pub/software/linux/tty

9.2 Installing chdrv

Unpack the binary distribution package

     # tar -xvzf chdrvbin-1.0.7.tar.gz
     # mv chdrvfont.tar.gz chdrv-1.0.7/
     # cd chdrv-1.0.7

Read the install instruction file INSTALL.1.0. Edit the script file installbin. If you are using shadow password for your system, you also need to edit the file chinese.conf, uncomment the line

LOGINPROGRAM /bin/telnet

and comment out the line

LOGINPROGRAM /bin/login

Now you are ready to go

     # ./installbin

9.3 GB Chinese Console - WZCE

Wei Zhong Chinese Environment is a software package support Chinese information processing on PC UNIX’s VGA console virtual screens.

9.3.1 Where is it

wzce_linux_elf_2.21n.tgz    ftp.ifcss.org:/pub/software/linux/tty

9.3.2 Installing

Read wzce linux.readme

10 Chinese WordProcessor – ChinesePower

A very user friendly WYSIWYG East Asia Word Processor on X window. Support GB, Big5, JIS and KS encoding mixed input, PS print, and display. It can generate gif file of the document with 7 colors. Cool for for generating Chinese Web Page. ChinesePower uses HBF or TTF fonts.

10.1 Where to get chpower

chpower-2.0.tar.gz       ftp.ifcss.org:/pub/software/x-win/editor/

10.2 Installing chpower

Chpower requires HBF fonts and also need Motif library to compile. First modify Makefile, then

     # make

will build the binaries for you. Also set some environment variables

     # setenv HBFPATH /usr/local/chinese/HBF/
     # setenv TTFPATH /usr/local/chinese/TTF/
     # setenv HZINPUTDIR /usr/local/chinese/dict/
     # setenv CHPOWERPATH path_of_chinesepower

11 Wabi 2.2 and Chinese window system

I have been using SoftPC on NextStep. I have Chinese Star intalled in SoftPC. It works great. But I have no extra bucks for Wabi 2.2, please let me know if you are using Chinese software in Wabi 2.2 for Linux.

12 Other Cool Stuffs

12.1 Crxvt

The xcin.tar.gz package includes a Chinese rxvt. It supports ANSI color and consumes less memory then cxterm. Xcin itself is an input server (Eten Input ?). But I use crxvt only

12.1.1 Getting Crxvt

xcin-2.1a.tgz       ftp.ifcss.org:/pub/software/x-win/xcin

12.1.2 Compiling and Installing

Here is the procedure to compile crxvt only (xcin may need ETEN system).

     # tar -xvzf xcin-2.1a.tgz
     # cd big5-pack/xcin
     # xmkmf
     # gcc -I/usr/X11R6/include -O2 -m486 -pipe -DDELETE_K=0 -DAuto_First=1 \
     > -c cli_xcin.c -o cli_xcin.o

     # cd ../crxvt
     # xmkmf
     # make
     # cp crxvt /usr/local/bin

12.2 Ktty

Ktty is another Hztty like utility for reading Chinese in kterm or pxvt. You can get ktty1.3.tar.gz from ftp.ifcss.org:/pub/software/unix/convert. This package does not compile on Linux. You need to use the file “tty.c” from the hztty-2.0.tar.gz package and also add two lines to the Makefile

linux:
      ${MAKE} ${MFLAGS} DEFINES='-DPOSIX -DSYSV -DUSE_SYSV_UTMP' ${PROG}

Now compile it by

     # make linux

12.3 TinTin++

TinTin++ is not a Chinese software. It is very useful when you are playing Chinese Mud. TinTin++ can be found at ftp.princeton.edu:/pub/tintin++. Here are three Chinese Mud sites

FengYun                  fengyun.com   5555
Xi You Ji                129.105.79.24 6905
Xia ke Xing              207.76.64.2   5555

13 Need more help ?

• rtfm.mit.edu:/pub/usenet/news.answers/chinese-text/big5-faq

• CXterm.FAQ

14 Appendix - Chinese Software Catalogue of FTP.IFCSS.ORG

Waring: The precompiled the packages for Linux are out of date, and may not working !

****
UNIX
****

========
UNIX:BBS
========
software: Phoenix BBS
version : 4.0, 1995.08.31
function: o UNIX platform multiuser BBS system,
          o discussion boards, private email
          o multichannel chat, one-to-one chat
          o Internet Email, News gateway,
          o 0Announce : Gopher-like information query interface.
          o fine tuned to allow more than 256 users on-line
          o configurable menu, screen display
          o Chinese message
URL     : ftp://ftp.ifcss.org/pub/software/unix/bbs/PhoenixBBS-4.0.tar.gz
author  : Ji-Tzay Yang, Ming-Feng Chen, Tzung-Yu Wen
============
UNIX:C-UTILS
============
software: addpy
version : 1.0
function: To annotate Pinyin to rare Hanzi and a portion of
          randomly selected common Hanzi.  Both GB (simplified) and
          Big5 (traditional) versions.  Based on statistics derived
          from huge Chinese corpus and well prepared hazni-freq-pinyin
          tables.  Source code and raw data provided.
URL     : ftp://ifcss.org/pub/software/unix/c-utils/addpy.tar.gz
filename: README.addpy, b5addpy.l, gbaddpy.l, b5addpy.dat,
          gbaddpy.dat, makefile
author  : Guo Jin
software: rm4mat
version : n/a
function: remove formating codes for printing in some GB files
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/rm4mat.c
author  : Chenghong Wang
software: mail_hxwz
version : 1.1
function: It extracts HXWZ from your mail (Suppose you are subscribing
          it).  It is a Bourne shell program which works on most Unix
          systems.  It can rerun itself every Friday automatically.

URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/mail_hxwz
author  : Yaoen Zhang
software: auto_get_hxwz
Version : 1.2
Function: It gets the current issue of HXWZ in GB or postscipt format.
          It also process and print these files, and do the clean up.
          It save your time and server machine time.
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/auto_get_hxwz
author  : Yaoen Zhang
software: Chinese less
version : 290, 1995.5.25
function: Browse BIG5 and GB text files
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/cless-290.tar.gz
author  : Paul W. Shew
comment : requires a chinese terminal, like cxterm.
software: cscreen
version : 3.2b
function: This is a modified screen to minic ETen in cxterm.
          After you run cscreen in cxterm, it will intercept
          every key movement and examine if the current
          screen postion has a Chinese char. If it has,
          auto-key movements will be made accordingly.
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/cscreen-3.2b.tar.gz
author  : Shih-Kun Huang
software: ytalk
version : 3.0.2c6
function: ytalk-3.02c provides better compatibility for DEC's
          workstations, two-byte refresh for terminal like cxterm, and
          capable of passing character code 254, internally used as erase
          function in ytalk-3.0; Compatible with ytalk-3.0, cytalk-3.0,
          cytalk-3.0.2, ytalk-3.02c0/1/2/3/4/5, that is, you can still use
          Delete, or Backspace key.  If both ends use ytalk-3.02c4, the code
          254 is transparent, you can use more Chinese words to talk to
          others.
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/ytalk-3.0.2c6.tar.gz
author  : Sze-Yao Ni
software: lunar
version : 2.1
function: conversion program between Solar and Chinese lunar calendars;

          calculation of birthday in "4-column" astrology as well
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/lunar-2.1.tar.gz
author  : Ricky Yeung and Fung Fung Lee
comment : can output GB code, bitmap, or just ASCII
software: DateStar - Chinese Calendar Producer
version : 1.1
function: Displays Chinese and western calendar in
          ASCII code, BIG-5 code (Hongkong, Taiwan),
              GuoBiao code (PRC Standard), and HZ code (Network)
          Prints on two most popular laser printers
              PostScript laser printers, and
              HP LaserJet (PCL) printers
          Generates four different format
              Yearly calendar, Monthly calendar,
              One-page Weekly calendar, and Two-page Weekly calendar
          Shows calendar with 24 Solar Terms (JieQi),
              the Heavenly Stems and Earthly Branches
              (TianGan DiZhi Eight Characters)
          Supports user defined annotations
          Applicable from year 1841 through to 2060
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/datestar-1.1.tar.gz
author  : Youzhen Cheng
comment : UNIX version works on SUN Workstation with SUN OS 4.1.x
          see /software/dos/c-utils for DOS version
software: pull
version : 2.9, 1996.4.7
function: 1. Extracts the original file(s) from uuencoded/compressed/split
             file(s).
             a) uudecode a file and display the name of its decoded file
             b) decompress .zz (= .gz/.Z/.zip) file
             c) uudecode + decompress
             d) uudecode + cat (concatenate) + gunzip
          2. Packs file. Reverse the above `extracting' process.
             compresses, (splits), then encodes file(s)
URL     : ftp://ftp.ifcss.org/pub/software/unix/utils/pull.c.gz
author  : Yao Li
software: utf-utils
version : 15 Oct 1993.
function: some utility programs for manipulating Unicode/ISO-10646 text
          in the FSS-UTF encoding.
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/utf-utils.tar.gz

author  : Ross Paterson
software: gbfmt
version : 1.0
function: GB formatting with variable line width, GB<->HZ
URL     : ftp://ftp.ifcss.org/pub/software/unix/c-utils/gbformat.tar.gz
author  : Dongxiao Yue (http://www.cs.umn.edu/~dyue/wiihist/gbfmt.html)
===========
UNIX:EDITOR
===========
software: celvis
version : 1.3
function: vi editor with Chinese handling capabilities.
          Most run on some Chinese terminal, e.g. cxterm, or IBM-PC with
          some kind of Chinese DOS.  Work exactly like Unix vi editor,
          except that it side-scrolls long lines instead of wrapping.
URL     : ftp://ftp.ifcss.org/pub/software/unix/editor/celvis.tar.gz
          ftp://ftp.cs.purdue.edu/pub/ygz
author  : Man-Chi Pong, Yongguang Zhang
comment : comments, bug-reports, modifications to: c-elvis@cs.purdue.edu
software: Chinese JOE (Joe's Own Editor)
version : 2.8c
function: Chinese BIG5 localized text screen editor.  The key sequences are
          remeniscent of WordStar and TURBO-PASCAL. Makes full use of
          termcap/terminfo, is designed to work well over slow networks and
          low baud rate modems, and has the best features of vi. Most
          useful for editing unformatted text, such as USENET news articles
          and for editing block-structured languages such as C and PASCAL.
URL     : ftp://ftp.ifcss.org/pub/software/unix/editor/joe2.8_c.tar.gz
author  : <u8222034@cc.nctu.edu.tw>
============
UNIX:CONVERT
============

software: cn2jp
version : 1.3.2, 1996.4.11
function: code conversion routines for Chinese and Japanese
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/cn2jp1.3.2.tar.gz
author  : Seke Wei

software: ktty
version : 1.3, 1996.4.11
function: This is a kanji terminal translator among Chinese and Japanese.
          It allows online translation of codes so that you can view a
          specific code real time using your Chinese or Japanese terminal.
          It runs on Unix and is derived from the 'hztty' package
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/ktty1.3.tar.gz
author  : Seke Wei
software: code1
version : 1.3, 1996.4.11
function: This is a multilingual file browser for Chinese and Japanese.
          It runs on Unix and allows browsing of files in various codes
          using a Chinese or Japanese terminal.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/code1.3.tar.gz
author  : Seke Wei
software: HUG program
version : 1.0, 1995.5.20
function: Converting between HZ, Uudecode/uuencode, GB codes
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/shug.osf.zip
          ftp://ftp.ifcss.org/pub/software/unix/convert/shug.sun4.zip
author  : Yinrong Huang
software: hztty
version : 2.0 Jan 29, 1994
Function: This program turns a tty session from one encoding to another.
          For example, running hztty on cxterm can allow you to read/write
          Chinese in HZ format, which was not supported by cxterm.
          If you have many applications in different encodings but your
          favor terminal program only supports one, hztty can make life easy.
          For example, hztty can your GB cxterm into a HZ terminal, a
          Unicode (16bit, or UTF8, or UTF7) terminal, or a Big5 terminal.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/hztty-2.0.tar.gz
author  : Yongguang Zhang
software: EHZ
version : 2.0
function: Conversion among GB/BIG5/CNS and EHZ-BIG5/GB/CNS.
          Patch to hztty to support EHZ-BIG5/GB/CNS.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/EHZ-2.0.tar.gz
author  : Ricky Yeung
comment : Also contains the EHZ spec and Fung Fung Lee's EHZ-BIG5-CNS spec.

software: b5cns
version : prototype
function: functions to convert Big5 <-> CNS
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/b5cns.tar.gz
author  : Ross Paterson
software: c2t
version : n/a
function: converts GB or BIG5 coded chinese to _pinyin_
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/c2t.tar.gz
authors : Tommi Kaikkonen and Katya Ta
software: BeTTY/CCF/B5Encode package
version : 1.534, 1995.03.22
function: a chinese code conversion package for codes widely used
          in Taiwan and the GB code widely used in Mainland, plus
          a 7-bit Big5 encoding method (B5Encode3/B5E3, an extension
          to HZ encoding for GB),
          including off-line converters (CCF/Chinese Code Filters and
          B5E/B5Encode) and an on-line converter (BeTTY) which simulates
          your native chinese terminal to become aware of the coding
          systems widely used in Taiwan and GB, HZ encoding.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/BeTTY-1.534.tar.gz
          ftp://hermes.ee.nthu.edu.tw/shin/betty/BeTTY-1.534.tar.gz
author  : Jing-Shin Chang
software: BeTTY-ws_2fl.p1
version : 1.0 (patch to BeTTY-1.534) Oct. 1995.
function: 1. makes BeTTY respect the window(tty) size.
          2. a second Chinese code converting filter can be added on.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/BeTTY-ws2fl.p1.tar.gz
author  : Wei Dong
software: c2gif
version : 0.01, 1995.10.21
function: convert a BIG5 text file to a GIF file
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/c2gif001.tar.gz
          http://www.math.ncu.edu.tw/~luors/c2gif/
author  : Luoh Ren-Shan
software: gb2jis
version : 1.5, 1995.11.19
function: convert GB (or HZ) to JIS with two-letter pinyin

URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/gb2jis.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/bdf/guobiao16.bdf.gz
author  : Koichi Yasuoka
software: jis2gb
version : 1.5, 1996.1.10
function: convert JIS to GB (or HZ)
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/jis2gb.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/bdf/guobiao16.bdf.gz
author  : Koichi Yasuoka
software: HZ
version : 2.0
function: convert among GB, HZ and zW
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/HZ-2.0.tar.gz
author  : Fung F. Lee
software: HZ+ specification and conversion utilities
version : 0.77
function: HZ+ is a convenient 7-bit representation of mixed Big5, GB,
          and ASCII text for use in Internet e-mail, news, etc.
          Source code for Big5 <-> HZ+ and GB <-> HZ+ conversion
          utilities is here.  DOS executables can be found in
          another archive, /software/dos/convert/hzp.zip.  New in
          this Unix version is a simple HZ+ terminal program for cxterm
          which allows the user to transparently read HZ+ mail and news.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/hzp.tar.gz
author  : Stephen G. Simpson
software: hc
version : 3.0
function: convert between GB and BIG5
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/hc-30.tar.gz
author  : Fung F. Lee and Ricky Yeung
software: Support Table for Hanzi Convert (hc)
version : 1994/05/01
function: Convert table supports the program Hanzi Convert
          (Author  : Fung F. Lee and Ricky Yeung) GB<->Big5.
          Include Russian, number, Japanese, graphing symbols and
          "incorrect" codes. Text file, comments in it.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/sym-supp.tab
          ftp://ftp.ifcss.org/pub/software/unix/convert/in-corr.tab
author  : Chi-Ming Tsai

software: pbmbig5
version : 0.01, 1995.11.2
function: convert big5 coded Chinese text file to pbm graphic file
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/pbmbig5-0.01.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/big/hbf/kck24.hbf
          ftp://ftp.ifcss.org/pub/software/fonts/big/hbf/kcchin24.f02
author  : Wei-Jou Chen
software: UTF utilities
version : 31 May 1994.
function: various utilities for the UTF encoding of Unicode/ISO-10646,
          including conversion from ISO-2022 and (partially) back again.
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/utf.tar.gz
author  : Ross Paterson
software: utf7
version : prototype, use at your own risk
function: functions to convert between UTF-7 and other codes
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/utf7.tar.gz
author  : Ross Paterson
software: ISO-2022-CN encoder and decoder
version : beta 960408
function: Convert between CN-GB and CN-CNS codes
URL     : ftp://ftp.ifcss.org/pub/software/unix/convert/iso-cn.tgz
author  : handa@etl.go.jp
comment : BIG5 code is not yet supported.
===============
UNIX:NETWORKING
===============
software: gopher2.014c
version : 2.0.14
function: A Chinese localized gopher client capable of 8-bit BIG5 Chinese
          string search on IBM AIX, SUN OS, and any other machines.
          Compatable with any BIG5 Chinese system such as ET and
          cxterm. executable binary for ibm and sun included.
URL     : ftp://ftp.ifcss.org/pub/software/unix/networking/gopher2.014c.tar.gz
          ftp://ftp.ifcss.org/pub/software/unix/networking/gopher.ibm.tar.gz
          ftp://ftp.ifcss.org/pub/software/unix/networking/gopher.sun.tar.gz
author  : Hoo-Tung Cheuk (NCTU CIS, Taiwan)

software: Chinese Tin
version : 1.2PL2a
function: tin 1.2PL2 newsreader with English/Big5 message toggle.
URL     : ftp://ftp.ifcss.org/pub/software/unix/networking/ctin122a.tar.gz
author  : Shih-Kun Huang
software: NcFTP with chinese message compatible
version : 2.3.0c, 1996.1.17
function: Chinese (BIG5) patch to NcFTP.
          Now it can display any chinese message from ftp server instead
          of "\xxx".  ANSI color compatible in Line Mode.
URL     : ftp://ftp.ifcss.org/pub/software/unix/networking/ncftp-2.3.0.chinese.tgz
author  : NCEMRSoft (orig), Aiken Sam (chinese patch)
==========
UNIX:PRINT
==========
software: C2PS
version : 1.30 Aug 1 1995
function: Translate Big5 coded Chinese document into Level 2 PostScript.
          This is the version for Sparcstation. Using Chinese TrueType fonts,
          you can create the most beautiful document with C2PS.
          This is a DEMO version. You can freely copied and use it.
URL     : ftp://ftp.ifcss.org/pub/software/unix/print/c2ps130sos.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/big5/ms-win/
author  : Hsueh-I Lu
software: cnprint
version : 2.60  JAN-25-95
function: print GB/Hz/BIG5/JIS/KSC/UTF8 etc or convert to PostScript
          (conforms to EPSF-3.0). Fast.  Multicolumn. Vertical printing.
          Small disk space requirement. "Intelligent" treatment of
          punctuations.  Flexible change of fonts, char size,
          width/height, char and line spaces, paper orientation and
          margins, etc.  Support of European chars.  Special modes for
          printing HXWZ.  See readme for more
URL     : ftp://ftp.ifcss.org/pub/software/unix/print/cnprint260.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/{gb,big5,misc,unicode}/hbf/
author  : Yidao Cai
comment : v2.60 is also for VMS, use v2.61 for DOS
software: GBscript
version : 1.11

function: Convert GB/ASCII mixed text to PostScript output.
          High print speed (4ppm on LaserWriter NTX).
          Support Adobe-2.1.
          Small PS file size (400K for one HXWZ issue).
URL     : ftp://ftp.ifcss.org/pub/software/unix/print/gbscript-1.11.tar.gz
author  : Yan Zhou
software: gb2ps
version : 2.02
function: convert GB/HZ to postscript, supports simple page formatting
          (change chinese fonts and font size, cover page, page
          number, etc). Five chinese fonts are provided in this
          release, they are Song, Kai, Fang Song, Hei and FanTi
          The HZ ENCODING is also supported.
URL     : ftp://ftp.ifcss.org/pub/software/unix/print/gb2ps.2.02.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/misc/fan24.ccf.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/misc/fang24.ccf.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/misc/hei24.ccf.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/misc/kai24.ccf.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/misc/song24.ccf.gz
author  : Wei SUN
software: news2ps
version : n/a
function: news2ps converts BIG5 to Postscript
URL     : ftp://ftp.ifcss.org/pub/software/unix/print/news2ps.c
          ftp://ftp.ifcss.org/pub/software/fonts/big5/misc/chinese.16.new
author  : The Society of HKU Postgraduate on Chinese Affairs
comment : rename chinese.16.new to chinese.16
software: hz2ps
version : 3.1
function: Convert hanzi (GB/BIG5) text to PostScript.
          Use HBF font files.
URL     : ftp://ftp.ifcss.org/pub/software/unix/print/hz2ps-3.1.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/{gb,big5}/hbf/
author  : Fung F. Lee
===========
UNIX:VIEWER
===========
software: ChiRK
version : 1.2a

function: GB/HZ/BIG5 text viewer on terminals (or emulations) capable
          of displaying Tektronics 401x graphics, such as GraphOn,DEC
          VT240/330, Xterm, Tektool on Sun, EM4105 on PC,
          VersaTerm-Pro on Mac, etc.
          displays up to 17x40 Chinese characters per screen. works directly
          under UNIX mail and news programs. Comes with four fonts.
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/ChiRK-1.2a.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/hbf/cclib.v
          ftp://ftp.ifcss.org/pub/software/fonts/gb/hbf/cclib.16
          ftp://ftp.ifcss.org/pub/software/fonts/gb/hbf/cclibf.16.gz
          ftp://ftp.ifcss.org/pub/software/fonts/big5/misc/chinese.16.new.gz
author  : Bo Yang
comment : rename chinese.16.new to chinese.16
software: Cbanner
version : 1.10, 950821
function: To show Chinese word's banner
URL     : ftp://ftp.ifcss.org/pub/software/unix/banner/cbanner1.10.tar.gz
author  : Sheen Cherng-Dar, rewritten by Jonen Liu
comment : requires ETen Big5 Chinese System's fonts.
software: gb2text
version : n/a
function: convert GB to text
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/gb2text.c
author  : Ding Yijun
software: hzbanner
version : 1.1, Feb 15, 1995
function: Display Song style GuoBiao in large ASCII characters,
          supports GB2312-80 (^[$A), GB2312-80 + GB8565-88 (^[$(E),
          Chinese-EUC (8-bit Guobiao) CNS Plane 1 & 2, BIG5 and HZ-encoding
          ( \c{c}·\~{a})
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/hzbanner11.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/gb/bdf/guobiao16.bdf.gz
          ftp://ftp.ifcss.org/pub/software/fonts/cns/bdf/cns1hku16.bdf.gz
          ftp://ftp.ifcss.org/pub/software/fonts/cns/bdf/cns2hku16.bdf.gz
author  : Koichi Yasuoka
software: hzview
version : 3.1
function: Display hanzi (GB/BIG5) text on dumb terminal.
          Use HBF font files.
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/hzview-3.1.tar.gz

          ftp://ftp.ifcss.org/pub/software/fonts/{gb,big5}/hbf/
author  : Fung F. Lee
Software: cnview
Version : 3.1 (UNIX version. DOS version available under /software/dos/viewer)
Function: View GB/Hz/Big5 encoded Chinese text file on * HP-UNIX (X-window) *
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/cnview.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/{big5,gb}/hbf/
Author  : Jifang Lin ({\ae}{\aa}?\`{e}^{3})
software: readgb
version : n/a
function: convert GB to text
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/readgb.c
author  : Yuzhao Lu
comment : modified from readnews.c
software: readnews
version : n/a
function: readnews converts BIG5 to ascii dot-matrix picture
URL     : ftp://ftp.ifcss.org/pub/software/unix/viewer/readnews.c
          ftp://ftp.ifcss.org/pub/software/fonts/big5/misc/chinese.16.new.gz
author  : The Society of HKU Postgraduate on Chinese Affairs
comment : rename chinese.16.new to chinese.16
==========
UNIX:INPUT
==========
software: CCTeach
version : 1.0
function: Chinese Character input method Teacher.
          Help new user to learn CC input and some utility programs
              of "cxterm NewFace" for associate dictionary tool,
              hotkey tool, WuBi phrase encoder, and converter with
              ".tit" <==> ".titnf".
          Based on GB and Big5 (ETen and HongKong).
          Support all input method by external dictionary file.
          Need cxterm in unix, CC DOS or ZW DOS in PC.
URL     : ftp://ftp.ifcss.org/pub/software/unix/input/CCTeach1.0.tar.gz
author  : Xiaokun Zhu
=========
UNIX:MISC

=========
software: ICCS 1.3
version : 1.3, June 26, 1994
function: Internet Chinese Chess Server
URL     : ifcss.org:/software/unix/misc/iccs-1.3.tar.gz
author  : Xi Chen
comment : file off-lined due to legal status July 3, 1996. Please contact
          the author Xi Chen at xichen@abel.math.harvard.edu for further info.
========
UNIX:WWW
========
software: cdelegate
version : 1.4a, 1996.5.1
function: This is a Chinese patch on DeleGate, a gateway for the WWW services.
          provides code translation between Chinese/Japanese for WWW browsers
URL     : ftp://ftp.ifcss.org/pub/software/unix/www/cdelegate1.4.tar.gz
author  : Seke Wei
software: Chinese Lynx
version : 2.5FMc, 1996.7.19
function: Chinese BIG5/GB patch to lynx, a WWW client for vt100 terminals.
          Volunteers needed to continue on the patch.
URL     : ftp://ftp.ifcss.org/pub/software/unix/www/clynx25.zip
author  : Nelson Chin
software: Internet MahJong Server (server + client applet)
version : 0.2beta
function: provides a server and a graphic client for playing MahJong on
          the Internet.
URL     : ftp://ftp.ifcss.org/pub/software/unix/www/MJ_dist.tar.gz
author  : Zuwei Thomas Feng
*****
LINUX
*****
software: D Series (Chinese Tools, ELF binary)
version : 1.00,  March 25, 1995
function: Binary distribution of various useful Chinese tools:
          Viewers -- cxterm, crxvt; Input server -- xcin; GB,Big5,HZ,B5E3
          code converters -- ccf, hc, hz2gb, gb2hz, zw2hz, b5decode,
          b5encode; Print tool -- lunar, ttf2ps; Pseudo tty -- hztty,

          betty; Various handy scripts, man pages, dictionaries, HBF fonts,
          xfonts included; chdrv, celvis, elm, sendmail
URL     : ftp://ftp.ifcss.org/pub/software/linux/CLinux/CTool/d1
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/CTool/d2
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/CTool/d3
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/CTool/d4
author  : Eric Lin
comment : requires XFree86 3.1+, ELF libraries
software: C Series (Chinese packages for Slackware)
version : N/A
function: The Chinese packages collected by Wei-Jou Chen can be installed by
          Slackware's setup tools. Basic idea are that we have right to install
          and remove Chinese softwares easily and beginners can play them
          without much trouble.
URL     : ftp://ftp.ifcss.org/pub/software/linux/CLinux/c1/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/c2/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/c3/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/c4/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/c5/
compiler: CHEN, Wei-Jou
software: MU Series (Mule packages for Slackware)
version : N/A
function: The Multilingual Emacs 2.0 packages for XFree86 2.X and 3.1
          collected by Shawn Hsiao can be installed by Slackware's setup
          tools
URL     : ftp://ftp.ifcss.org/pub/software/linux/CLinux/mu1/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/mu2/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/mu3/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/mu4/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/mu5/
          ftp://ftp.ifcss.org/pub/software/linux/CLinux/mu6/
compiler: Shawn Hsiao
=============
LINUX:C-UTILS
=============
software: GNU fileutils-3.9
version : 3.9, 1 August 1994
function: Chinese version of the GNU file utility programs for Linux.
URL     : ftp://ftp.ifcss.org/pub/software/linux/c-utils/fileutils-3.9-bin-chinese.tar.gz
author  : Patrick D'Cruze

software: C2PS
version : 1.30 Aug 1 1995
function: Translate Big5 coded Chinese document into Level 2 PostScript.
          This is the version for Linux. Using Chinese TrueType fonts, you
          can create the most beautiful Chinese documents with C2PS.
          This is a DEMO version. You can freely copied and use it.
URL     : ftp://ftp.ifcss.org/pub/software/linux/chinese_utils/c2ps130lnx.tar.gz
          ftp://ftp.ifcss.org/pub/software/fonts/big5/ms-win/
author  : Hsueh-I Lu
============
LINUX:EDITOR
============
software: \~\{a\}\{\ae\}\frac{1}{4}?\{\ae\}¸ \~\{a\}for Linux (promotion version)
version : v2.163
function: PE2-like text editor, special designed for Chinese
URL     : ftp://ftp.ifcss.org/pub/software/
agent   : LU, Heman ({\aa}{\ae}?·\'{e})
=========
LINUX:TTY
=========
software: chdrv
version : 1.0.7, 1995.12.20
function: Chinese Terminal Simulator. Does not require X-Windows.
URL     : ftp://ftp.ifcss.org/pub/software/linux/tty/chdrv-1.0.7.tar.gz
          ftp://ftp.ifcss.org/pub/software/linux/tty/chdrvbin-1.0.7.tar.gz
          ftp://ftp.ifcss.org/pub/software/linux/tty/chdrvfont.tar.gz
author  : WANG, Yu-Chung
===========
LINUX:X11R6
===========
software: Behavior DTop (for Linux) [{\aa}?{\ae}¸ \"{a}¸ ?\`{e}±{\ae}{\ae}\c{c}\c{c}^{3}>>\c{c}?±]
version : 1.4, Beta
function: A full-featured Chinese DeskTop Publishing Software Package
          characterized by object-oriented design for manipulating
          various document objects, including text, tables, graphics,
          equations, images, in an integrated way. Two outline fonts
          are provided in the Beta Version. PostScript output.

          Good as an English DeskTop Publishing Software Package too.
          (See the README's & formated DTop manual files for a long and
          complete list of functions and characteristics).
URL     : ftp://ftp.ifcss.org/pub/software/linux/X11R6/dtop1.4/
author  : DTop Development Group